OpenStudy (jennychan12):

Find inverse derivative of f(x) at point a. when f(x) = tan x when a = sqrt 3 /3

OpenStudy (jennychan12):

Find \[f^{-1}(x)\] when \[a = \frac{ \sqrt 3 }{ 3 }\] \[f(x) = \tan x\]

OpenStudy (jennychan12):

i'm really confused on this question.

OpenStudy (anonymous):

\[y=\tan^{-1}(x)\]

OpenStudy (jennychan12):

i haven't learned derivative of arctan

OpenStudy (anonymous):

yes i will prove to you the derivative ...

OpenStudy (jennychan12):

huh? i just need to solve the question. i don't really need a proof -_- would you take the derivative of sec x?

zepdrix (zepdrix):

Hmm I'm not really familiar with this type of problem. I had to look it up on Wikipedia. This is the formula I'm getting for the derivative of the inverse function.\[\large \left[f^{-1}\right]'(a)=\frac{1}{f'\left(f^{-1}(a)\right)}\] So I think what we need to do issss, let's start by finding \(f^{-1}(a)\).

zepdrix (zepdrix):

\[\large f(x)=\tan x \qquad \rightarrow \qquad f^{-1}(x)=\arctan x\]

OpenStudy (anonymous):

well you can use the triangle method for this or use implicit differentiation

OpenStudy (jennychan12):

@mathsmind ...what? so then x = tan y so x = pi/6 right?

zepdrix (zepdrix):

Our \(a\) value is given to be \(\dfrac{\sqrt3}{3}\).\[\large f^{-1}\left(\dfrac{\sqrt3}{3}\right)=\arctan \left(\frac{\sqrt3}{3}\right)\] Pi/6? Hmm yes that sounds right! :) good.

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

but you need to find the derivative

OpenStudy (jennychan12):

er yeah. i kinda skipped a step. would you take the derivative of the original equation? so f'(x) = sec^2 x and just plug in pi/6 and then flip it??

zepdrix (zepdrix):

I think that brings us to this, \[\large \left[f^{-1}\right]'\left(\frac{\sqrt3}{3}\right)=\frac{1}{f'\left(\dfrac{\pi}{6}\right)}\] I can stop with the fancy notation if it's confusing you :) lol

zepdrix (zepdrix):

Yes that sounds right :)

OpenStudy (jennychan12):

okay. cuz i have the work for the answers but i was confused on what they did. thanks :)

zepdrix (zepdrix):

cool c: